# Can LLMs Beat Humans in Debating? A Dynamic Multi-agent Framework for Competitive Debate

### v1

@@ Abstract
Competitive debate is a comprehensive and complex computational argumentation task. Large Language Models (LLMs) encounter hallucinations and lack competitiveness in this task. To address these challenges, we introduce Agent for Debate (Agent4Debate), a dynamic, multi-agent framework based on LLMs designed to enhance their capabilities in competitive debate. Drawing inspiration from human behavior in debate preparation and execution, Agent4Debate employs a collaborative architecture where four specialized agents (Searcher, Analyzer, Writer, and Reviewer) dynamically interact and cooperate. These agents work throughout the debate process, covering multiple stages from initial research and argument formulation to rebuttal and summary. To comprehensively evaluate framework performance, we construct the Chinese Debate Arena, comprising 66 carefully selected Chinese debate motions. We recruite ten experienced human debaters and collect records of 200 debates involving Agent4Debate, baseline models, and humans. The evaluation employs the Debatrix automatic scoring system and professional human reviewers based on the established Debatrix-Elo and Human-Elo ranking. Experimental results indicate that the state-of-the-art Agent4Debate exhibits capabilities comparable to those of humans. Furthermore, ablation studies demonstrate the effectiveness of each component in the agent structure.
@@ Introduction
In recent years, Large Language Models (LLMs) (OpenAI 2023; Touvron et al. 2023b) have demonstrated remarkable capabilities in various natural language processing tasks, offering new possibilities for constructing high-performance debate systems. However, in the specific context of competitive debate, LLMs still need to overcome two significant challenges. First, LLMs often face hallucination problems (Ji et al. 2023), where models may generate plausible information that is inaccurate or fabricated. Second, due to limitations in safety alignment during training (Ouyang et al. 2022) and constraints in handling long contexts (Liu et al. 2024), models often need to improve in adversarial and sustained debate scenarios (shown in Figure 1), struggling to maintain competitiveness and argumentative consistency.
@@
To address these challenges, we propose a multi-agent framework based on LLMs, Agent for Debate (Agent4Debate). Agent4Debate features a dynamic, multi-agent collaborative architecture, leveraging the cooperation of multiple specialized models to enable the framework to participate in multi-stage competitive debates. Our framework demonstrates performance comparable to the human-level in competitive debate. To comprehensively evaluate the competitive debate capabilities of Agent4Debate, we develop the Chinese Debate Arena. This arena comprises 66 carefully selected Chinese debate motions, covering three categories (Abell 2018), such as Policy, Value, and Fact, thoroughly testing the performance of participants across different types of debates. Participants include Agent4Debate with different foundation models, two baselines, and ten experienced human debaters. All participants engage in pairwise matches, with each debate assessed through two independent evaluation methods, including an automatic debate judging system based on the Debatrix (Liang et al. 2024)metrics, and an expert judging system consisting of three human reviewers. Based on these two sets of independent evaluation results, we construct two separate Elo (Elo 1967; Zheng et al. 2023) ranking lists, providing a multi-faceted quantitative assessment of participants’ performance across various debate motions. The experimental results from the arena demonstrate that Agent4Debate can achieve human-level performance in various types of competitive debates, as evidenced by Debatrix and human judgments.
@@
  • We propose the Agent4Debate, which enhances the performance of LLMs in competitive debates through dynamic multi-agent collaboration. This framework mimics human debate team interactions, with agents adapting roles and strategies.
  • We construct the Chinese Debate Arena, comprising 66 carefully selected Chinese debate motions and 200 debate matches across Policy, Value, and Fact categories. Human debaters are incorporated, and we establish Debatrix-Elo and Human-Elo rankings using Debatrix metrics and professional human judges, respectively. Results indicate that Agent4Debate’s performance in competitive debates is comparable to that of humans. Ablation studies validate the effectiveness of each component within this flexible, human-inspired framework.
@@
Argumentation research has deep historical roots (Walton, Reed, and Macagno 2008), with its core objective being to achieve persuasion through logical reasoning and promote consensus among parties. In recent years, computational argumentation has emerged as an increasingly important field in natural language processing, with its main research directions encompassing argument mining (Lawrence and Reed 2019; Chen et al. 2024), argument generation (Hua, Hu, and Wang 2019), argument persuasiveness (Carlile et al. 2018), and argument quality assessment (Wachsmuth et al. 2017b; Liang et al. 2024; Wachsmuth et al. 2024). With the rise of Large Language Models (LLMs), research utilizing adversarial methods such as debate to enhance model capabilities (Du et al. 2023; Chang 2024) has gradually attracted academic attention. Against this backdrop, our study focuses on competitive debate, a complex computational argumentation task that integrates multiple sub-tasks.
@@
LLMs, such as ChatGPT (OpenAI 2023), LLaMA (Touvron et al. 2023b, a), demonstrate powerful capabilities in instruction following and reasoning tasks. Harnessing these advanced capabilities, researchers have developed LLM-based agents, which mark a significant step forward in the field. These agents leverage the language understanding and generation abilities of models for more sophisticated tasks like multi-step reasoning and interactive problem-solving, as shown in recent studies (Wang et al. 2023b; Li et al. 2023). They find uses across various domains, such as software engineering (Qian et al. 2023) and scientific inquiry (Boiko, MacKnight, and Gomes 2023), highlighting their versatility. These agents can imitate complex human actions, partake in social interactions (Park et al. 2023; Tu et al. 2023), and replicate intricate scenarios like elections (Argyle et al. 2022), debates (Wang et al. 2023a; Du et al. 2023), and consumer patterns (Wang et al. 2023c), illustrating their capacity to emulate human social dynamics.
@@
Figure 2: Competitive debate format.
Competitive debate is a structured multi-turn interactive task. Each turn of statement can be regarded as a document-level text generation task, with a temporal and logical progression relationship between multiple turns. A typical debate has two opposing sides: the Pro side and the Con side. We represent the competitive debate as an interleaved sequence:
@@
where $(s_{i},r_{i})$ denotes the $i$-th statement and its corresponding role, $s_{i}$ is the statement, and $r_{i}\in\{\text{Pro},\text{Con}\}$ represents the role of speaker. Each statement can be defined as:
$s_{i}=\mathcal{G}(m,r_{i},D_{(i-1)})$ (2)
where $m$ is the motion of debate, $D_{(i-1)}$ represents the history of the first $i-1$ statements, and $\mathcal{G}(\cdot)$ is the generation function that produces each statement.
Our debate structure comprises three distinct stages, such as constructive arguments, rebuttals, and summary statements. The format is illustrated in Figure 2. To ensure fairness and simulate actual competitive debate conditions (Whitman 2005), we establish specific rules for each stage:
@@
The collaboration in Agent4Debate is not just a simple sequence of steps, but rather a dynamic interaction between multiple agents, based on the debate stage and context. All agents are equipped with customized prompts for different debate stages, enabling them to better adapt to and execute the specific tasks of the current stage. In the following sections, we introduce the functions of each agent in detail.
@@
The Searcher is a tool agent in the Agent4Debate framework, designed to effectively mitigate hallucination issues and address information timeliness problems that LLMs may encounter during debates. It achieves this by accessing and organizing information from external knowledge bases. The workflow of Searcher primarily involves decomposing search questions into more refined queries, then utilizing external tools (such as search engines or specialized knowledge bases) to retrieve relevant information, and finally systematically compiling and organizing the obtained answers. The information compiled by the Searcher forms a static motion knowledge base, which is accessible to all agents for reference throughout the entire debate process. This approach ensures consistency and reliability of the information used in the debate. Note that, the Searcher plays slightly different roles at various stages of the debate. In Stage 1, the Searcher uses the motion as the search question for information gathering. However, in Stage 2 and Stage 3, the Searcher switches to a passive mode, waiting for specific instructions from the Writer before conducting targeted searches.
@@
The Analyzer is a core agent in the Agent4Debate framework, integrating real-time information from the debate and providing structured guidance for subsequent content output. Its primary function is to systematically analyze and plan the debate content based on the given motion, current stage, and historical context, thus bridging different phases of the debate. The workflow of Analyzer primarily involves breaking down the debate content step-by-step, drafting detailed outlines, and providing targeted strategic advice to other agents. This approach ensures coherence in debate reasoning and comprehensiveness in argumentation. Notably, the Analyzer plays slightly different roles at various stages of the debate:
@@
The Reviewer maintains argumentative coherence by continuously assessing consistency with previously presented information across all debate stages. This process involves providing feedback and modification suggestions to the Writer, facilitating targeted revisions. The review-revision cycle persists iteratively until the script meets the Reviewer’s quality standards.
@@
Our experimental design involves three types of subjects, such as the baseline framework, Agent4Debate based on different LLMs, and human participants. For all models, we set temperature to 0.2 and Top P to 0.75, with no other parameters adjusted.
@@
We adopt the benchmark framework of AI-Debater 2024 competition  [^1] http://www.fudan-disc.com/sharedtask/AIDebater24, incorporating Tavily [^2] https://tavily.com as the search engine and stage-specific prompts. We uses Claude-3.5-sonnet and Deepseek-Chat as the foundation model.
@@
To comprehensively evaluate the generalization capability of Agent4Debate and conduct more in-depth comparative experiments, we select a variety of advanced LLMs as the foundation for Agent4Debate. These models include Claude-3.5-sonnet, GPT-4o (OpenAI 2023), and Gemini-1.5-Pro/Flash (Reid et al. 2024), all of which have demonstrated excellent performance in various evaluations (Zheng et al. 2023). Considering that our study focuses on Chinese competitive debate, we specifically incorporate several LLMs that excel in Chinese language processing, including Qwen2-72b-Instruct (Yang et al. 2024), Deepseek-Chat-v2 (Bi et al. 2024), and GLM-4-Air (Zeng et al. 2022). Switching models in Agent4Debate experiments updates all components accordingly. In all experiments, the searcher used Tavily as the search engine.
@@
We recruit ten experienced debaters for our experiment to validate the performance of Agent4Debate against humans in competitive debate. Each with 2-4 years of debate team training and at least one year of Chinese competitive debate experience. These debaters are not involved in other aspects of this study’s development. They are informed that they will be debating against artificial intelligence and are given 1-3 days of preparation time for each motion. To ensure effective communication, we use the Whisper model (Radford et al. 2023) to transcribe human speeches into text while the human debaters read the model’s output directly. This design ensured accurate information transfer and provided human debaters ample time for reflection and response.
@@
To comprehensively assess the abilities of Agent4Debate, Baseline, and Humans in competitive debate, we establish the Competitive Debate Arena. This arena is designed to provide a comprehensive and fair evaluation environment, covering various types of debate motions and assessment methods. We carefully select 66 debate motions from major Chinese debate competitions over the past decade, including Chinese Debate World Cup, The World Mandarin Debating Championship, and International Chinese Debating Competition. These motions cover three main categories (Abell 2018): Value, Fact, and Policy. Fact makes statements or comparisons about testable aspects of the natural world, Value assigns value or judgment to certain things or concepts, while Policy typically suggests action plans through proposed changes.
@@
This weight function adjusts the importance of each match in the final ranking based on score differences, making the ranking calculation more precise. Based on this weight function, our likelihood function becomes:
$L(\gamma)=\prod_{i=1}^{n}P(A_{i}>B_{i})^{w_{i}}$ (5)
@@
Our improved Elo system not only effectively reflects participants’ overall performance in multiple matchups but also allows for more nuanced adjustments based on the specifics of each match. Using two independent review methods and ranking systems, we can better understand the performance of participants and compare potential differences between Debatrix and human reviews.
@@
Baseline Comparison Study
We conduct a comparative performance evaluation of Agent4Debate against the baseline. Each framework participates in 20 debates across five different motions. To ensure fairness, the number of times each framework argued for the Pro and Con sides is balanced. Debatrix are employed as the evaluation criteria. Debatrix scoring is applied three times for each debate, with 1 point awarded for each win in the dimensions of Argument (A), Language (L), Source (S), and Overall (O) performance. In the case of a tie, both sides are awarded 0.5 points.
@@
Table 1: The results of comparison experiment.
As shown in Table 1, Agent4Debate enhances the competitive debating performance across both models. For Claude-3.5-sonnet, the Overall score improves from 0.38 to 2.62, while for Deepseek-Chat, it increases from 0.23 to 2.77. These results demonstrate that the Agent4Debate framework effectively enhances the performance of language models of varying scales and types in competitive debate tasks. Among all metrics, Source shows the most significant improvement. This can be attributed to the Searcher Agent and Analyzer Agent within Agent4Debate, which conducts an in-depth analysis of debate motions and systematic organization of materials, utilizing external knowledge more effectively than the simple search approach from baseline. The Language metric shows relatively modest improvement, reflecting robust generation capabilities of LLMs, leaving limited room for enhancement.
Comparing the results between Claude-3.5-sonnet and Deepseek-Chat, it is observed that Agent4Debate yields more pronounced performance improvements for larger models, particularly in the Argument and Overall metrics. This may be due to larger models possessing more vital reasoning abilities and better instruction-following capabilities (Kaplan et al. 2020), thus exhibiting superior adaptability to complex frameworks.
@@
Results of Chinese Debate Arena
To comprehensively evaluate debate performance, we conduct a large-scale experimental assessment. We collect records of 200 debate matches (excluding those from comparison experiments and ablation studies), covering 66 debate motions across three categories, including Fact, Value, and Policy. Participants included Agent4Debate using different foundation models, two baselines, and ten human debaters, all of whom engaged in randomly paired competitions. Each debate is independently assessed using both the Debatrix and human judges. Utilizing the improved BT model introduced earlier, we calculate Elo scores for all 200 matches and sub-Elo scores for each of the three debate categories. The experimental results are presented in two independent ranking systems: Debatrix-Elo (Table 3) and Human-Elo (Table 4). Models without an asterisk (*) indicate the foundation models used by Agent4Debate, while those with an asterisk denote the models used as baselines.
@@
Table 3: The results of Debatrix-Elo Ranking.
@@
Table 4: The results of Human-Elo Ranking.
Agent4Debate, especially those using advanced foundation models such as Gemini-1.5-Pro and Claude-3.5-sonnet, demonstrate performance comparable to or surpassing human debaters in both Debatrix-Elo and Human-Elo rankings. The top-performing Agent4Debate (Gemini-1.5-Pro) consistently ranks first, scoring 1044.18 in Debatrix-Elo and 1040.64 in Human-Elo. Experimental results indicate that models with more robust reasoning and instruction-following capabilities perform better within the Agent4Debate framework.
In Debatrix-Elo, most models show score variations across the Fact, Policy, and Value categories. In contrast, Human-Elo displays more consistent scores for each model across categories. This disparity may arise because Debatrix considers Source, Language, and Argument dimensions, while human judges likely focus more on logic and rebuttal techniques. Debatrix-Elo and Human-Elo show high consistency in model rankings, particularly for top-performing models. However, human performance is ranked differently in the two rankings. In Debatrix-Elo, humans rank 8th with a score of 978.35, while in Human-Elo, they rank 4th with a score of 1006.46. This suggests that Debatrix-Elo may underestimate human performance. This underestimation is partly due to the different evaluation tendencies between Debatrix and human judges, and partly because human speech quality deteriorates when transcribed to text.
In Debatrix-Elo, certain models excel in specific categories. This is due to significant differences in the argumentation processes for the three types of debate motions: Policy debates typically require extensive evidence to demonstrate policy necessity and effectiveness; Value debates often demand more substantial logical reasoning and expressive skills; Fact debates combine characteristics of both. These distinctions, reflected in Debatrix’s multi-dimensional evaluation, yield varying results.
Detailed Performance Analysis
We conduct a separate analysis of 30 debates between Agent4Debate and human debaters. In these debates, to ensure comprehensive experimentation, all foundation models of Agent4Debate have participated. The scoring results from the Debatrix system and human judges are presented in Table 5.
@@
Table 5: Comparison of Human and Agent4Debate Performance in Chinese Debate Arena
Debatrix for human performance is lower than human judges across three dimensions. This discrepancy may stem from several factors. Regarding Source, human debaters use voice input, which is then transcribed into text. People typically do not directly cite references in oral debates, leading to lower scores. The Language score is the lowest, possibly due to oral expressions often containing verbal tics and informal language, coupled with imperfect voice-to-text transcription accuracy, affecting language quality assessment. The low Argument score may be a cascading effect of the previous two low scores, thus impacting Debatrix’s overall understanding and evaluation of human input.
In contrast, human judges employ different criteria when evaluating competitive debates. They usually prioritize core factors such as logical reasoning and debating skills, only considering other aspects when these primary elements are challenging to distinguish. This approach to judgment differs significantly from the Debatrix.
Table 6 presents the consistency results between humans and the Debatrix. The results show that internal consistency among human reviewers remains stable across all matches, while the consistency between Debatrix and human reviewers varies when including or excluding human debaters. These findings further corroborate the above observations. In this analysis, tie is considered to be a consistent outcome.
@@
Table 6: Consistency between Debatrix and Human Judges in Chinese Debate Arena
Although Debatrix shows considerable differences from human reviewers in evaluating debates between humans and models, this does not imply that Debatrix is an entirely unreliable indicator for assessing competitive debates. Particularly in evaluating debates between models, Debatrix can provide multi-faceted analytical results, which are still valuable for analyzing the comprehensive capabilities of models.

### v2

@@ Abstract
Competitive debate is a complex task of computational argumentation. Large Language Models (LLMs) suffer from hallucinations and lack competitiveness in this field. To address these challenges, we introduce Agent for Debate (Agent4Debate), a dynamic multi-agent framework based on LLMs designed to enhance their capabilities in competitive debate. Drawing inspiration from human behavior in debate preparation and execution, Agent4Debate employs a collaborative architecture where four specialized agents, involving Searcher, Analyzer, Writer, and Reviewer, dynamically interact and cooperate. These agents work throughout the debate process, covering multiple stages from initial research and argument formulation to rebuttal and summary. To comprehensively evaluate framework performance, we construct the Competitive Debate Arena, comprising 66 carefully selected Chinese debate motions. We recruit ten experienced human debaters and collect records of 200 debates involving Agent4Debate, baseline models, and humans. The evaluation employs the Debatrix automatic scoring system and professional human reviewers based on the established Debatrix-Elo and Human-Elo ranking. Experimental results indicate that the state-of-the-art Agent4Debate exhibits capabilities comparable to those of humans. Furthermore, ablation studies demonstrate the effectiveness of each component in the agent structure.
@@ Introduction
In recent years, Large Language Models (LLMs) (OpenAI 2023; Touvron et al. 2023b) have demonstrated remarkable capabilities in various natural language processing tasks, offering new possibilities for constructing high-performance debate systems. Competitive debate, characterized by multi-turn document-level text generation with inter-turn logical dependencies, presents a unique challenge for LLMs, particularly in two significant areas. First, LLMs often face hallucination problems (Ji et al. 2023), where models may generate plausible information that is inaccurate or fabricated. Second, due to limitations in safety alignment during training (Ouyang et al. 2022) and constraints in handling long contexts (Liu et al. 2024), models often need to improve in adversarial and sustained debate scenarios (shown in Figure 1), struggling to maintain competitiveness and argumentative consistency.
@@
To address these challenges, we propose a multi-agent framework based on LLMs, Agent for Debate (Agent4Debate). Agent4Debate features a dynamic, multi-agent collaborative architecture, leveraging the cooperation of multiple specialized LLMs to enable the framework to participate in multi-stage competitive debates. Inspired by human debate preparation processes, our framework incorporates four key agents, including Searcher, Analyzer, Writer, and Reviewer. To comprehensively evaluate the competitive debate capabilities of Agent4Debate, we establish the Competitive Debate Arena, employing an Elo ranking system widely used in competitive sports, ensuring fairness and scalability. This arena comprises 66 carefully selected Chinese debate motions, covering three categories (Abell 2018), such as Policy, Value, and Fact, thoroughly testing the performance of participants across different types of debates. Participants include Agent4Debate with different foundation models, two baselines, and ten experienced human debaters. All participants engage in pairwise matches, with each debate assessed through two independent evaluation methods, including an automatic debate judging system based on the Debatrix (Liang et al. 2024) metrics, and an expert judging system consisting of three professional human reviewers. Based on these two sets of independent evaluation results, we construct two separate Elo (Elo 1967; Zheng et al. 2023) ranking lists, providing a multi-faceted quantitative assessment of participants’ performance across various debate motions. The experimental results from the arena demonstrate that Agent4Debate can achieve human-level performance in various types of competitive debates.
@@
  • We propose the Agent4Debate, which enhances the performance of LLMs in competitive debates through dynamic multi-agent collaboration. This framework mimics human debate team interactions, with agents adapting roles and strategies. Specifically, it employs the Searcher for information gathering, the Analyzer for strategic assessment, the Writer for argument formulation, and the Reviewer for critical evaluation.
  • We construct the Competitive Debate Arena, a public resource comprising 66 Chinese debate motions and 200 debate matches across Policy, Value, and Fact categories. Human debaters are incorporated, and we establish Debatrix-Elo and Human-Elo rankings using Debatrix metrics and professional human judges, respectively.
  • Our experimental results indicate that Agent4Debate’s performance in competitive debates is comparable to that of humans. Ablation studies validate the effectiveness of each component.
@@
Argumentation research has deep historical roots (Walton, Reed, and Macagno 2008), with its core objective being to achieve persuasion through logical reasoning and promote consensus among parties. In recent years, computational argumentation has emerged as an increasingly important field in natural language processing, with its main research directions encompassing argument mining (Lawrence and Reed 2019; Chen et al. 2024), argument generation (Hua, Hu, and Wang 2019), argument persuasiveness (Carlile et al. 2018), and argument quality assessment (Wachsmuth et al. 2017b; Liang et al. 2024; Wachsmuth et al. 2024). Project Debater (Slonim et al. 2021), a debate system that integrates multiple modules, relies on retrieval-based methods rather than generative approaches for its argumentation. With the rise of Large Language Models research utilizing adversarial methods such as debate to enhance model capabilities (Du et al. 2023; Chang 2024) has gradually attracted academic attention. Against this backdrop, our study focuses on competitive debate, a complex computational argumentation task that integrates multiple sub-tasks.
@@
LLMs, such as ChatGPT (OpenAI 2023), LLaMA (Touvron et al. 2023b, a), demonstrate powerful capabilities in instruction following and reasoning tasks. Harnessing these advanced capabilities, researchers have developed LLM-based agents, which mark a significant step forward in the field. These agents leverage the language understanding and generation abilities of models for more sophisticated tasks like multi-step reasoning and interactive problem-solving, as shown in recent studies (Wang et al. 2023b; Li et al. 2023). They have various uses across different domains, such as software engineering (Qian et al. 2023) and scientific inquiry (Boiko, MacKnight, and Gomes 2023), highlighting their versatility. These agents can imitate complex human actions, partake in social interactions (Park et al. 2023; Tu et al. 2023), and replicate intricate scenarios like elections (Argyle et al. 2022), debates (Wang et al. 2023a; Du et al. 2023), and consumer patterns (Wang et al. 2023c), illustrating their capacity to emulate human social dynamics. While these agents demonstrate impressive capabilities in emulating human social dynamics, current research predominantly explores collaborative scenarios. However, competitive settings, though equally crucial in human interactions, remain comparatively underexplored.
@@
Competitive debate is a structured multi-turn interactive task. Each turn of statement can be regarded as a document-level text generation task, with a temporal and logical progression relationship between multiple turns. A typical debate has two opposing sides: the Pro side and the Con side. We represent the competitive debate as a sequence:
@@
where $(s_{i},r_{i})$ denotes the $i$-th statement and its corresponding role, $s_{i}$ is the statement, and $r_{i}\in\{\textit{Pro},\textit{Con}\}$ represents the role of speaker. Each statement can be defined as:
$s_{i}=\mathcal{G}(m,r_{i},D_{i-1})$ (2)
where $m$ is the motion of debate, $D_{i-1}$ represents the history of the first $i-1$ statements, and $\mathcal{G}(\cdot)$ is the generation function that produces each statement.
Typical competitive debate structure usually comprises three distinct stages, namely constructive arguments, rebuttals, and summary statements. To ensure fairness and simulate actual competitive debate conditions (Whitman 2005), we establish specific rules for each stage:
@@
The collaboration in Agent4Debate is not just a simple sequence of steps, but rather a dynamic interaction between multiple agents, based on the debate stage and context. All the agents are equipped with customized prompts for different debate stages, enabling them to better adapt to and execute the specific tasks of the current stage. In the following sections, we introduce the functions of each agent in detail.
@@
The Searcher is a tool agent in the Agent4Debate framework, designed to effectively mitigate hallucination issues and address information timeliness problems that LLMs may encounter during debates. It achieves this by accessing and organizing information from external knowledge bases. The workflow of Searcher primarily involves decomposing search questions into more refined queries, then utilizing external tools (such as search engines or specialized knowledge bases) to retrieve relevant information, and finally systematically compiling and organizing the obtained answers. The information compiled by the Searcher forms a motion knowledge base, which is fixed and accessible to all agents for reference throughout the entire debate process. This approach ensures consistency and reliability of the information used in the debate. Note that, the Searcher plays different roles at various stages of the debate. In Stage 1, the Searcher uses the motion as the search question for information gathering. However, in Stage 2 and Stage 3, the Searcher switches to a passive mode, waiting for specific instructions from the Writer before conducting targeted searches.
@@
The Analyzer is a core agent in the Agent4Debate framework, integrating real-time information from the debate and providing structured guidance for subsequent content output. Its primary function is to systematically analyze and plan the debate content based on the given motion, current stage, and historical context, thus bridging different phases of the debate. The workflow of Analyzer primarily involves breaking down the debate content step-by-step, drafting detailed outlines, and providing targeted strategic advice to other agents. This approach ensures coherence in debate reasoning and comprehensiveness in argumentation. Notably, the Analyzer plays different roles at various stages:
@@
The Reviewer maintains argumentative coherence by continuously assessing consistency with previously presented information across all debate stages. This process involves providing feedback and modification suggestions to the Writer, and facilitating targeted revisions. The review-revision cycle persists iteratively until the script meets the Reviewer’s quality standards.
@@
Our experimental design involves three types of participants, including the baseline framework, Agent4Debate based on different LLMs, and human participants. For all models, we set temperature to 0.2 and Top P to 0.75, with no other parameters adjusted.
@@
We adopt the benchmark framework of AI-Debater 2024 competition [^1] http://www.fudan-disc.com/sharedtask/AIDebater24, incorporating Tavily [^2] https://tavily.com as the search engine and stage-specific prompts. We uses Claude-3.5-sonnet and Deepseek-Chat (Bi et al. 2024) as the foundation models.
@@
To comprehensively evaluate the generalization capability of Agent4Debate and conduct more in-depth comparative experiments, we select a variety of advanced LLMs as the foundation for Agent4Debate. These models include Claude-3.5-sonnet, GPT-4o (OpenAI 2023), and Gemini-1.5-Pro/Flash (Reid et al. 2024), all of which have demonstrated excellent performance in various evaluations (Zheng et al. 2023). Considering that our study focuses on Chinese competitive debate, we specifically incorporate several LLMs that excel in Chinese language processing, including Qwen2-72b-Instruct (Yang et al. 2024), Deepseek-Chat , and GLM-4-Air. Switching models in Agent4Debate experiments updates all components accordingly. In all experiments, the Searcher used Tavily as the search engine.
@@
We recruit ten experienced debaters for our experiment to validate the performance of Agent4Debate against humans in competitive debate. Each debater has with 2-4 years of debate team training and at least one year of Chinese competitive debate experience. They are informed that they will be debating against artificial intelligence and are given 2 days of preparation time for each motion. To ensure effective communication, we use the Whisper model (Radford et al. 2023) to transcribe human speeches into text while the human debaters read the model’s output directly. This design ensures accurate information transfer and provides human debaters ample time for reflection and response. These debaters participate only in the debates, not in other research activities.
@@
To comprehensively assess the abilities of Agent4Debate, Baseline, and Humans in competitive debate, we establish the Competitive Debate Arena. This arena is designed to provide a fair and extensible evaluation environment, covering various types of debate motions and assessment methods. We carefully select 66 debate motions from major Chinese debate competitions over the past decade, including Chinese Debate World Cup, The World Mandarin Debating Championship, and International Chinese Debating Competition. These motions cover three main categories (Abell 2018), including Value, Fact, and Policy. Fact makes statements or comparisons about testable aspects of the natural world, Value assigns value or judgment to certain things or concepts, while Policy typically suggests action plans through proposed changes.
@@
where $\text{score}\in[0,3]$. This weight function adjusts the importance of each match in the final ranking, making the ranking calculation more precise. Based on this weight function, our likelihood function becomes:
$\mathcal{L}=\prod_{i=1}^{n}P(A_{i}>B_{i})^{w_{i}}$ (5)
@@
Our improved Elo system not only effectively reflects participants’ overall performance in multiple matchups but also allows for more nuanced adjustments based on the specifics of each match. Using two independent review methods and ranking systems, we can better understand the performance of participants and compare potential differences between Debatrix and human reviews. Furthermore, this Elo system is scalable, efficiently incorporating new frameworks or models for ongoing comparative analysis.
@@
Agent4Debate vs. Baseline
We conduct a comparative performance evaluation of Agent4Debate against the baselines. Each framework participates in 20 debate matches, including five different motions. To ensure fairness, the number of times each framework argued for the Pro and Con sides is balanced. Debatrix is employed as the evaluation criteria. Debatrix scoring is applied three times for each debate, with 1 point awarded for each win in the dimensions of Argument (A), Language (L), Source (S), and Overall (O) performance. In the case of a tie, both sides are awarded 0.5 points.
@@
Table 1: Comparison of Agent4Debate and Baseline.
As shown in Table 1, Agent4Debate enhances the competitive debating performance across both models. For Claude-3.5-sonnet, the Overall score improves from 0.38 to 2.62, while for Deepseek-Chat, it increases from 0.23 to 2.77. These results demonstrate that the Agent4Debate framework effectively enhances the performance of language models of varying scales and types in competitive debate tasks. Among all metrics, Source shows improvement. This can be attributed to the Searcher Agent and Analyzer Agent within Agent4Debate, which conducts an in-depth analysis of debate motions and systematic organization of materials, utilizing external knowledge more effectively than the simple search approach from baseline. The Language shows relatively modest improvement, reflecting robust generation capabilities of LLMs, leaving limited room for enhancement.
Comparing the results between Claude-3.5-sonnet and Deepseek-Chat, it is observed that Agent4Debate yields more pronounced performance improvements for the more powerful model, particularly in the Argument and Overall metrics. This may be due to more advanced models possessing stronger reasoning abilities and better instruction-following capabilities (Kaplan et al. 2020), thus exhibiting superior adaptability to complex frameworks.
@@
Results of Competitive Debate Arena
We collect records of 200 debate matches (excluding those from comparison experiments and ablation studies), covering 66 debate motions across three categories, including Fact, Value, and Policy. Participants included in Agent4Debate using different foundation models, two baselines, and ten human debaters, all of which are engaged in randomly paired competitions. Each debate is independently assessed using both the Debatrix and human judges. Utilizing the improved BT model in Eq. 5, we calculate Elo scores for all 200 matches and sub-Elo scores for each of the three debate categories. The experimental results are presented in two independent ranking systems, consisting of Debatrix-Elo (Table 3) and Human-Elo (Table 4).
@@
Table 3: Debatrix-Elo Ranking. denotes baseline models, unmarked models are Agent4Debate foundation models.
@@
Table 4: Human-Elo Ranking. denotes baseline models, unmarked models are Agent4Debate foundation models.
Drawing from the experimental results presented in Tables 3 and 4, we can derive the following insights. (1) Agent4Debate, especially those using advanced foundation models such as Gemini-1.5-Pro and Claude-3.5-sonnet, demonstrate performance comparable to or surpassing human debaters in Debatrix-Elo and Human-Elo rankings. The top-performing Agent4Debate (Gemini-1.5-Pro) consistently ranks first, scoring 1044.18 in Debatrix-Elo and 1040.64 in Human-Elo. Experimental results indicate that models with more robust reasoning and instruction-following capabilities perform better within the Agent4Debate framework. (2) In Debatrix-Elo, most models show score variations across the Fact, Policy, and Value categories. In contrast, Human-Elo displays more consistent scores for each model across categories. This disparity may arise because Debatrix considers Source, Language, and Argument dimensions, while human judges likely focus more on logic and rebuttal techniques. Debatrix-Elo and Human-Elo show high consistency in model rankings, particularly for top-performing models. However, human performance is ranked differently in the two rankings. In Debatrix-Elo, humans rank 8th with a score of 978.35, while in Human-Elo, they rank 4th with a score of 1006.46. This suggests that Debatrix-Elo may underestimate human performance. This underestimation is partly due to the different evaluation tendencies between Debatrix and human judges, and partly because human speech quality deteriorates when transcribed to text. (3) In Debatrix-Elo, certain models excel in specific categories. This is due to differences in the argumentation processes for the three types of debate motions: Policy debates typically require extensive evidence to demonstrate policy necessity and effectiveness; Value debates often demand more substantial logical reasoning and expressive skills; Fact debates combine characteristics of both. These distinctions, reflected in Debatrix’s multi-dimensional evaluation, yield varying results.
Agent4Debate vs. Humans
@@
Table 5: Comparison of Human and Agent4Debate
We conduct a separate analysis of 30 debates between Agent4Debate and human debaters. In these debates, to ensure comprehensive experimentation, all foundation models of Agent4Debate participate. The scoring results from the Debatrix system and human judges are presented in Table 5.
Debatrix for human performance is lower than human judges across three dimensions. This discrepancy may stem from several factors. Regarding Source, human debaters use voice input, which is then transcribed into text. People typically do not directly cite references in oral debates, leading to lower scores. The Language score is the lowest, possibly due to oral expressions often containing verbal tics and informal language, coupled with imperfect voice-to-text transcription accuracy, affecting language quality assessment. The low Argument score may be a cascading effect of the previous two low scores, thus impacting Debatrix’s overall understanding and evaluation of human input. In contrast, human judges employ different criteria when evaluating competitive debates. They usually prioritize core factors such as logical reasoning and debating skills, only considering other aspects when these primary elements are challenging to distinguish. This approach to judgment differs from the Debatrix.
Consistency
To analyze the differences between Debatrix and human evaluations, we conduct a consistency analysis. Consistency is calculated by comparing the result between human and Debatrix, with tie considered consistent outcomes. Table 6 presents the results, showing that internal consistency among human reviewers remains stable across all matches, while the consistency between Debatrix and human reviewers varies when including or excluding human debaters. These findings further corroborate the above observations. These findings suggest that while Debatrix shows differences from human reviewers in evaluating debates between humans and models, it still provides valuable insights, particularly in assessing model-to-model debates. In these cases, Debatrix offers multi-faceted analytical results that contribute to our understanding of models’ comprehensive capabilities in competitive debates.
@@
Table 6: Consistency between Debatrix and Human Judges
We further analyze the Elo rankings and Agent4Debate in the Appendix.